OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (anonymous):

Every sequence is divided by 3 every time: So, 24/3 = 8 8/3 = 2.66 2.66/3 = 0.886..

OpenStudy (anonymous):

Thank you!